OpenStudy (anonymous):

The Choo family is putting new tiles in their entryway .each tile is a 1 foot squeare. if the area of entrywaty is 28 square feet which of the following could be the perimeter of entryway? A. 20 ft. B. 22 ft. C. 26 ft. D. 28 ft.

Directrix (directrix):

The perimeter of a rectangle is the sum of the distances around it. The rectangle in this problem is the entry way with area 28. To get the area, you do length times width. You need the length and width to get the perimeter. So, look at the factors of 28. See what they add up to and check against options in the problem.

Directrix (directrix):

28 = 28 times 1 28 = 2 times 14 28 = 4 times 7

Directrix (directrix):

So, on 28 = 28 times 1, the length is 28 and the width is 1. Perimeter = 2 lengths and 2 widths added. Two lengths = 56 and 2 widths is 2 for a sum of 58. There is no option of 58. So go to next set of factors.

Directrix (directrix):

28 = 2 times 14 That means a length of 2 and a width of 14. For perimeter, that would be two times 2 and then two times 14 = 4 + 28 = 32. There is no 32 option so go to next set of factors.

OpenStudy (anonymous):

Wow! Thank u!! ^-^

Directrix (directrix):

28 = 4 times 7 means a length of 4 and a width of 7. Two of the 4s and 2 of the 7s = 8 + 14 = 22 for perimeter. Is there an option of 22? Yes, so B) 22 is the perimeter.

Directrix (directrix):

Question?